OpenStudy (anonymous):

Which of the following is a polynomial with roots:

OpenStudy (anonymous):

\[2, -2i , 2i\]

OpenStudy (anonymous):

the one that looks like \[(x-2)(x+2i)(x-2i)\]

OpenStudy (anonymous):

or what you get when you multiply this out, which is not as hard as it seems since \[(x+2i)(x-2i)=x^2+4\]so your real job is to multiply \[(x-2)(x^2+4)\]

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

except.. these are what I can choose from for this question: x3 – 4x2 + 4x – 16 x3 + 4x2 + 4x + 16 x3 + 2x2 + 4x + 8 x3 – 2x2 + 4x – 8

OpenStudy (anonymous):

there is no -5 in the problem multiply \[(x-2)(x^2+4)\] and see which one you get

OpenStudy (anonymous):

..I got -5?..

OpenStudy (anonymous):

\[(x-2)(x^2+4)\] is a polynomial, not a number

OpenStudy (anonymous):

i.e. multiply out using the distributive law. sometimes known by the unfortunate acronym "foil"

OpenStudy (anonymous):

im confused. I do not understand

OpenStudy (anonymous):

your job is to find a polynomial with three zeros, namely \(2,-2i,2i\) the way you can find this polynomial is first write it in factored form, as \[(x-2)(x+2i)(x-2i)\] just like if you were going to solve \(x^2-3x+2=0\) you would write in factored form as \((x-1)(x-2)=0\) in this case you know the zeros, so you know how to factor your job here is to multiply out just like if you multiply \((x-1)(x-2)\) you get \((x-1)(x-2)=x^2-3x-2\)

OpenStudy (anonymous):

therefore what you need to do to solve is this problem is to multiply three factors together you must multiply \((x-2)(x+2i)(x-2i)\)

OpenStudy (anonymous):

first you multiply the two terms \((x+2i)(x-2i) =x^2+4\) and then you multiply \[(x-2)(x^2+4)\] and see which if the three answers above you get

OpenStudy (anonymous):

it is clear how to multiply these two terms together?